OpenStudy (anonymous):

given sin theta=4/5 with 90 dgreess less than or equal to thata less than or equal to 180 degress find exact value of : sin theta/2

OpenStudy (anonymous):

\[\sin \theta=\ 4/5 \ with 90 degreess\] =\[180 degreess \] so its 1:2

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u=(2 cos pi/4)i + (2 sin pi/4)j, v = (cos 3pi/2)i + (sin 3pi/2)j find the angle theta .... If sin theta equals two third and zero degrees less than theta less than 90

OpenStudy (anonymous):

ok tnaks

OpenStudy (anonymous):

yes i need help

OpenStudy (kropot72):

sin theta = 4/5 = 0.8 sin^-1 theta = 53.13010235 degrees sin (53.13010235/2) = 0.447213595

OpenStudy (anonymous):

no but not indecimals

OpenStudy (anonymous):

do it in a graph

OpenStudy (kropot72):

I have the answer: sin (theta/2) = Square root (1/5) Do you want to know the method?

OpenStudy (kropot72):

The length of the remaining side of the triangle is: square root of (5^2 - 4^2) = 3 Then use the formula for the sine of the half angle: sin (theta/2) = square root (((s - b)(s - c))/bc)) Where s = 1/2(a + b + c)